archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    CLOSED Percent Complete Details An incident has been detected on SFR equipment in the departments 60 and 80 we will investigate Date Tuesday 29 October 2013 20 06PM Reason for closing Done Comments 1 Comment by OVH Tuesday 29 October

    Original URL path: http://status.ovh.es/?do=details&id=5686 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    cut LNS Bordeaux Closed 2012 11 29 3725 ADSL SDSL Maintenance Backend Core FS 7667 Marseille SHE13 Closed 2012 11 22 3727 ADSL SDSL Modernization Backend Core FS 7669 Update Switchs Closed 2012 11 22 3722 ADSL SDSL Maintenance Backend Core FS 7662 Routing infra unbuilding Paris Closed 2012 11 21 3720 ADSL SDSL Maintenance Backend Core FS 7663 NRA VAU75 Closed 2012 11 21 3709 ADSL SDSL Modernization Backend Core FS 7651 lil 3 4 n5 Closed 2012 11 20 3695 ADSL SDSL Maintenance Backend Core FS 7635 Dep 35 Collect Closed 2012 11 15 3680 ADSL SDSL Incident Backend Core FS 7622 Modems configuration Closed 2012 11 14 3647 ADSL SDSL Maintenance Backend Core FS 7589 ADVA Equipement on SHE13 Closed 2012 11 07 3628 ADSL SDSL Incident Backend Core FS 7571 Collect in 22 Closed 2012 11 05 3590 ADSL SDSL Modernization Backend Core FS 7533 Upgrade switchs Closed 2012 10 30 3587 ADSL SDSL Maintenance Backend Core FS 7530 hsrp fix on th2 11 n5 and gsw 11 n5 Closed 2012 10 28 3579 ADSL SDSL Maintenance Backend Core FS 7522 Collection in the 70 Closed 2012 10 26 3552 ADSL SDSL Maintenance Backend Core FS 7494 Optical loop Paris RUE92 VIR78 CAC94 VEL78 Closed 2012 10 18 3543 ADSL SDSL Modernization Backend Core FS 7485 switches Paris Closed 2012 10 17 3524 ADSL SDSL Maintenance Backend Core FS 7466 lil 3 n5 lil 4 n5 Closed 2012 10 17 3539 ADSL SDSL Maintenance Backend Core FS 7481 Update of can59 gob75 and vau75 Closed 2012 10 16 3533 ADSL SDSL Maintenance Backend Core FS 7475 Updating was59 Closed 2012 10 16 3516 ADSL SDSL Incident Backend Core FS 7458 Collecte Dep 30 Closed 2012 10 12 3513 ADSL SDSL Incident Backend Core FS 7455 NRA

    Original URL path: http://status.ovh.es/?pagenum=6&string=&type=&sev=&dev=&due=&cat=&status=&order2=lastedit&sort=asc&sort2=desc&date=&project=&order=proj (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Type Incident Category Backend Core Status CLOSED Percent Complete Details We are having an incident on the optical loop in Bordeaux The dslam aqu33 is now accessible Date Tuesday 22 October 2013 16 11PM Reason for closing Done Additional comments about closing We have replaced the faulty device The service is back up since 13 00 Comments 1 Comment by OVH Tuesday 22 October 2013 09 51AM We have detected

    Original URL path: http://status.ovh.es/?do=details&id=5641 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    There is an issue on the SFR collect in the South of France A ticket has been opened the investigations are underway on their side Date Monday 21 October 2013 00 45AM Reason for closing Done Comments 2 Comment by OVH Friday 18 October 2013 16 34PM The fault has been identified on the Bordeaux BAS SFR has just switched onto an emergency BAS The connections are going back up

    Original URL path: http://status.ovh.es/?do=details&id=5621 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Core Status CLOSED Percent Complete Details There is an issue in départements 25 and 70 on the SFR collect Investigations are underway on their side Date Wednesday 16 October 2013 16 40PM Reason for closing Done Comments 1 Comment by

    Original URL path: http://status.ovh.es/?do=details&id=5598 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Task Type Modernization Category Backend Core Status CLOSED Percent Complete Details We will intervene on Paris optical loop to insert 10G lambdas and to improve the security of the collect France Telecom No impact on traffic is expected The operation will be performed Thursday October 10th starting from 11am Customers of collecte France Telecom Ile de France will be switched tonight to collecte Lille to avoid having desynchronization during the

    Original URL path: http://status.ovh.es/?do=details&id=5568 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Status CLOSED Percent Complete Details Tuesday 1st October at Midnight we will activate VDSL2 for our eligible customers connected to Main Distribution Frame divided by OVH A downtime of a few minutes is planned during the resynchronisation of the modem in VDSL2 Date Tuesday 01 October 2013 08 58AM Reason for closing Done Comments 2 Comment by OVH Tuesday 01 October 2013 08 57AM We started switching for eligible customers

    Original URL path: http://status.ovh.es/?do=details&id=5504 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Date Friday 27 September 2013 16 52PM Reason for closing Done Comments 5 Comment by OVH Friday 27 September 2013 12 09PM A team in currently on site Comment by OVH Friday 27 September 2013 12 38PM We found a cut on the 7km fibre 3km from the MDF The fibre maintenance team will arrive at 13 30 Comment by OVH Friday 27 September 2013 15 03PM The teams are

    Original URL path: http://status.ovh.es/?do=details&id=5495 (2015-10-15)
    Open archived version from archive



  •